UT March 18, 2019

OSA: D.H.Huerta

ISp: J. Morris

ISA; B. Rothberg (Tucson)

Mtn Manager: P. Hartley

ARGOS AO: S. Walsh (Tucson)

ARGOS Commandant: G. Rahmer (Tucson)

Observers: B. Rothberg, S. Walsh, G. Rahmer (Tucson); AZ 2nd half - SO: S. Walsh (Tucson)

UT 01:25 (and earlier) - A. Cavallaro - AO FLAO checkout of WFS - completed

UT 01:29 - LUCI Software - started up - took two darks L2 luci2.20190318.0001-0002

UT 01:49 - P. Hartley check DIMM for water marks/damage - moving to Horizon

UT 02:32 - rotating LUCI-2 to 210 - will run init - taking two more test darks 0003-0004

UT 03:02 - sending ARGOS collimation script to M67

-WFS sensor looks....odd...triangular PSF spots are not illuminated uniformly

-too much trefoil, astigmatism? Z10 is 1600, Z6 is 4200...

UT 03:22 - S Walsh and G. Rahmer working on lasers

UT 03:23 - I cleared active optics - see if this helps with collimation

-that seemed to help a lot. Z10 is ~500

UT 03:41 -sending science preset

UT 03:44 - looks like it found the wrong AO Ref (or the script was the faint one - AO Ref should be 12.4R, but pyramid is reporting 14.4)

0005 - bad image

- was not able to abort 2nd image in series

- the abort just prevented it from saving

UT 03:46 - reseding script

Pointed too close to the Moon

UT 03:48 -moving to NGC 2782

UT 04:17 - still difficulties with Laser system - currently collimating on star - but collimation looks terrible - z10 runaway again

UT 04:22 - lost guidestar

UT 04:32 - looks like collimation is back under control - but motor issues plaguing the laser system

UT 04:37 - powering down the laser system - will restart to check whether lasers can move

UT 04:50 - collimation went bad (we weren't doing much anyway) - active optcs being cleared

UT 05:06 - restart of systems - still issue moving a motor switch with one of the lasers - going to fire on-sky to see which way the lasers do or don't move - could explain wobbling

-GCS crashed

UT 05:34 - ready to test

UT 05:37 - LUCI-2 rotator limit was 270seconds away....slewed to Park Position to avoid issue.

UT 05:39 - resending preset to NGC 2782 (collimation preset)

-no guidestar found, but should be able to run ARGOS test

-satellite closure

UT 05:58 - testing laser - moving around manually to see if it matches he jitter seen previously

UT 06:02 - looks like ARGOS is done for the night - motor control issue for the lasers seems to be a show stopper - will reconfigure soon for AZ service observing (MODS-Bino)

UT 06:13 - reconfiguring for Bino-MODS

-- BarryRothberg - 18 Mar 2019
Topic revision: r1 - 18 Mar 2019, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback